# Tonal coarticulation revisited: functional covariance analysis to investigate the planning of co-articulated tones by Standard Chinese speakers

### v1

@@ 1 Introduction
GAMMs are a powerful and flexible tool, but they also rely on a ”discretized” version of the f0 curves, which are instead usually measured in Hz and most often assumed to be smooth and continuous throughout their trajectories. The intrinsic smoothness of the f0 curves points naturally to the functional data analysis (FDA) domain (Ramsay and Silverman [22], Ferraty and Vieu [7]) as a tool to study them. Accounting for the smooth structure of the underlying process helps in dealing with the high dimensionality of the data objects. Previous works which applied functional data analysis to linguistic data included lip-motion Ramsay et al. [23], analysis of temporal prosodic effects Byrd, Krivokapić and Lee [2] and speech production Koenig, Lucero and Perlman [11]. Recently however there seems to be a surge of successful application of FDA to linguistic, as it became apparent from the recent works Tavakoli and Panaretos [25], Aston, Chiou and Evans [1], Pigoli et al. [21], Koshy and Tavakoli [12], Gubian, Torreira and Boves [8], who use various aspect of functional data, from functional dimentionality reduction analysis in Gubian, Torreira and Boves [8] to second order analysis via functional covariances in Aston, Chiou and Evans [1], Pigoli et al. [21].
@@
The novelty of our work consists in proposing a unified approach to study tonal variation of f0 curves both with respect to their shape, as well as to their variation of second order, and to apply it to a dataset specifically collected to study the effect of memory task on coarticulation. To the best of our knowledge, no study considered tonal coarticulation as a second-order effect. As a first exploratory step, we will investigate the effect of cognitive load on vowel duration via a Cox regression model (Cox [6]). Effect of cognitive load on f0 contour is then explored via GAMMs. In the particular dataset considered, generalized additive models do not seem to provide a conclusive evidence on the cognitive load effect on coarticulation, in a way disproving empirical evidence by experts while at the same time suggesting that if such effect is significant, it might be found at finer resolution. Thus, we moved onto analysing the covariance structure of the f0 residuals curves, obtained after the effect on the mean of each tonal combination was modelled via the GAMM model. We employed the functional ANOVA permutation test by Masarotto, Panaretos and Zemel [16], which was showed to have state of the art performance, to show that indeed, the cognitive load effect is a higher resolution effect.
@@
The article is structured as follows. Section 2 describes the data collection procedure and results based on [33]. The statistical tools employed are described in Section 3. In particular, duration is investigated with survival methods as described in Section 3.1 and the procedure behind the first order analysis via GAMMs can be found in Section 3.2. Section 3.3 lays out the necessary definitions in functional data analysis, Section 3.4 describes a functional test that allows to test the hypothesis of equality between covariance operators with very high power, and Section 3.5 describes the procedure behind the only instances of tangent space principal component analysis known to investigate modes of variation among covariance operators. Results are collected in Section 4.
@@ 3 Statistical methods
3.1 Analysis of the duration data
Following Xu [32], the duration of the first and second vowel in the /mama/ sequences were tested. A Cox proportional-hazard regression model Cox [6] was estimated to study the association between cognitive load and and vowel duration. Time fixed covariates include the presence of cognitive load, the different tone and the interactions between such terms. Every recording is treated as a singular observation, and recordings are clustered by speakers. Separate analysis for f0 contours were built for each target tone (first and second syllable separately). All analyses were performed in R software environment (https://www.R-project.org/) with the mstate library.
@@ 4 Results
This Section will first present the study over the durational patterns of the tone-bearing syllables, followed by first and second order analysis the interaction effects of tonal context and cognitive load. In particular, functional ANOVA allows to detect a significant difference between the covariance operators relative to curves with cognitive load and without cognitive load, as presented in Section 4.3. Covariance principal component is used to tap into such difference in Section 4.4.
4.1 Duration of the tone-carrying vowels
[…]
Effect on first syllable HR $95\%$ CI p-value
Cognitive Load 1.01 $[0.90,1.16]$ $0.82$
First syllable T2 0.79 $[0.69,0.91]$ $0.05^{*}$
First syllable T3 0.79 $[0.69,0.90]$ $0.05^{*}$
First syllable T4 1.03 $[0.93,1.16]$ $0.53$
Second syllable T2 0.98 $[0.88,1.11]$ $0.78$
Second syllable T3 0.88 $[0.70,1.11]$ $0.26$
Second syllable T4 1.03 $[0.95,1.12]$ $0.44$
Effect on second syllable HR $95\%$ CI p-value
Cognitive Load 1.21 $[0.99,1.45]$ $0.05^{*}$
First syllable T2 1.00 $[0.84,1.20]$ $0.98$
First syllable T3 0.96 $[0.81,1.14]$ $0.62$
First syllable T4 0.97 $[0.84,1.14]$ $0.78$
Second syllable T2 0.72 $[0.52,1.02]$ $0.05^{*}$
Second syllable T3 0.59 $[0.41,0.87]$ $<0.01^{*}$
Second syllable T4 4.59 $[3.22,6.56]$ $<0.01^{*}$
[…]
@@ 5 Discussions
We would like to emphasize that more follow-up analysis are possible, for example, once an effect of cognitive load at covariance level is detected, pairwise comparison to tap into this effect in the fashion of Pesarin and Salmaso [20] is also possible. We leave the details to future work.

### v2

@@ 1 Introduction
GAMMs are a powerful and flexible tool, but they also rely on a ”discretized” version of the f0 curves, which are instead usually measured in Hz and most often assumed to be smooth and continuous throughout their trajectories. The intrinsic smoothness of the f0 curves points naturally to the functional data analysis (FDA) domain (Ramsay and Silverman [22], Ferraty and Vieu [7]) as a tool to study them.
Accounting for the smooth structure of the underlying process helps in dealing with the high dimensionality of the data objects. Previous works which applied functional data analysis to linguistic data included lip-motion Ramsay et al. [23], analysis of temporal prosodic effects Byrd, Krivokapić and Lee [2] and speech production Koenig, Lucero and Perlman [11]. Recently however there seems to be a surge of successful application of FDA to linguistic, as it became apparent from the recent works Tavakoli and Panaretos [25], Aston, Chiou and Evans [1], Pigoli et al. [21], Koshy and Tavakoli [12], Gubian, Torreira and Boves [8], who use various aspect of functional data, from functional dimentionality reduction analysis in Gubian, Torreira and Boves [8] to second order analysis via functional covariances in Aston, Chiou and Evans [1], Pigoli et al. [21].
@@
The novelty of our work consists in proposing a unified approach to study tonal variation of f0 curves both with respect to their shape, as well as to their variation of second order, and to apply it to a dataset specifically collected to study the effect of memory task on coarticulation. To the best of our knowledge, no study considered tonal coarticulation as a second-order effect. Effect of cognitive load on f0 contour is then explored via GAMMs. In the particular dataset considered, generalized additive models do not seem to provide a conclusive evidence on the cognitive load effect on coarticulation, in a way disproving empirical evidence by experts while at the same time suggesting that if such effect is significant, it might be found at finer resolution. Thus, we moved onto analysing the covariance structure of the f0 residuals curves, obtained after the effect on the mean of each tonal combination was modelled via the GAMM model. We employed the functional ANOVA permutation test by Masarotto, Panaretos and Zemel [16], which was showed to have state of the art performance, to show that indeed, the cognitive load effect is a higher resolution effect.
@@
The article is structured as follows. Section 2 describes the data collection procedure and results based on [33]. The statistical tools employed are described in Section 3. In particular, the procedure behind the first order analysis via GAMMs can be found in Section 3.1. Section 3.2 lays out the necessary definitions in functional data analysis, Section 3.3 describes a functional test that allows to test the hypothesis of equality between covariance operators with very high power, and Section 3.4 describes the procedure behind the only instances of tangent space principal component analysis known to investigate modes of variation among covariance operators. Results are collected in Section 4.
@@ 4 Results
This Section presents the first and second order analysis the interaction effects of tonal context and cognitive load. In particular, functional ANOVA allows to detect a significant difference between the covariance operators relative to curves with cognitive load and without cognitive load, as presented in Section 4.2. Covariance principal component is used to tap into such difference in Section 4.3.
@@ 5 Discussions
We would like to emphasize that more follow-up analysis are possible, for example, once an effect of cognitive load at covariance level is detected, pairwise comparison to tap into this effect in the fashion of Pesarin and Salmaso [20] is also possible. Moreover, one could investigate the effect of cognitive load on vowel duration via a Cox regression model (Cox [6]). In this setting, time fixed covariates include the presence of cognitive load, the different tone and the interactions between such terms. We leave the details to future work.